19. You lend a friend $10,000 for which your friend will repay you $27,027 at the end of 5 years. What interest rate are you only questions 19 and 21. please show all steps
Answer #1

19.

Calculating the Interest Charged using the Interest Compounded Formula:-

Future Value = Principal (1+Interest rate)No of Period

where, Future Value = $ 27027

Principal = $10,000

No of Period = 5

Note- We are assuming that interest is componded annually.

27027 = 10000(1+r)5

2.7027 = (1+r)5

taking 5root on both sides,

1.22 = (1+r)

r = 1.22-1

r = 0.22 or 22%

So, Interest rate charged is 22% compounded annually.

Note- As HOMEWORKLIB Guidelines, I'm only allowed to answer 1 question per answer so Question-19 has been answered.

If you need any clarification regarding this solution, then you can ask in comments

If you like my answer then please Up-vote as it will be motivating.